OpenStudy (anonymous):

A light bulb manufacturer randomly tests 1500 bulbs and finds that 2 are defective. How many defective bulbs would be expected in a shipment of 24,000 bulbs? A. 32 B. 48 C. 120 D. 750

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

\[\frac{ 2 }{ 1500 }=\frac{ x }{ 24000 }\] Cross multiply and get x alone

OpenStudy (anonymous):

48000/1500x

OpenStudy (anonymous):

1500x=48000 x=48000/1500

OpenStudy (anonymous):

320 lightbulbs are defective

OpenStudy (anonymous):

Are you sure? I got 32

OpenStudy (anonymous):

thats what i meant lol i have been typing in so many numbers today its not even funny

OpenStudy (anonymous):

Either way, that would be the answer.

OpenStudy (anonymous):

lol yeah